A Stone has an initial velocity of 200 ft per see up to the right at a slope of 4 to 3. The components of acceleration are constant at a_x=-12 ft per second² and a_y=-20 ft per second². Compute the radius of curvature at the start and at the top of the path



1
Expert's answer
2021-04-01T01:04:32-0400
"a_x=-12, a_y=-20"

"\vec r''(t)=-12\vec i-20\vec i"

"\vec v_0=200\cdot\dfrac{3}{\sqrt{3^2+4^2}}\vec i+200\cdot\dfrac{4}{\sqrt{3^2+4^2}}\vec j"

"\vec v_0=120\vec i+160\vec j"

"\vec r'(t)=(120-12t)\vec i+(160-20t)\vec j"


"\vec r'(t)\times \vec r''(t)=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 120-12t & 160-20t & 0 \\
 -12t & -20t & 0 
\end{vmatrix}"

"=\vec k((120-12t)(-20t)-(160-20t)(-12t))"

"=-480t\vec k"

"||\vec r'(t)\times \vec r''(t)||=480|t|"

"||\vec r'(t)||=\sqrt{(120-12t)^2+(160-20t)^2}"


"=4\sqrt{34t^2-580t+2500}"

"K(t)=\dfrac{||\vec r'(t)\times \vec r''(t)||}{||\vec r'(t)||^3}"

"=\dfrac{480t}{64(34t^2-580t+2500)^{3/2}}"

"=\dfrac{15t}{2(34t^2-580t+2500)^{3/2}}"

"R(t)=\dfrac{2(34t^2-580t+2500)^{3/2}}{15t}"

At the start of the path: "t=0"


"R(0)=\infin"

At the top of the path: "v_y=0"


"160-20t=0"

"t=8"

"R(8)=\dfrac{2(34(8)^2-580(8)+2500)^{3/2}}{15(8)}=3.6 (ft)"